# Unsupervised Anomaly Detection through Mass Repulsing Optimal Transport

### v1

@@ 3.3 Building and Generalizing an Anomaly Score
Through MROT, we want to build a score for samples, based on how anomalous they are. Assuming some samples lie in a low density region of space, our MROT problem forces those samples to send their mass to distant parts of the feature space. In contrast, normal samples can send their mass to the immediate neighborhood outside the exclusion zone. As a result, we can sort out normal from anomalous samples using the transportation effort,
@@
$\mathcal{T}_{i}=\sum_{j=1}^{n}\dfrac{\gamma_{ij}^{\star}}{p_{i}}\tilde{c}_{k}(x_{i}^{(P)},x_{j}^{(P)}),$ (9)
@@
where $p_{i}$ is the importance of the $i-$th sample (e.g, $p_{i}=n^{-1}$ for uniform importance). Interestingly, equation 9 is similar to the barycentric map, widely used in domain adaptation Courty et al. (2016); Montesuma et al. (2023). $\mathcal{T}$ has 2 shortcomings as an anomaly score. First, it is hardly interpretable, as its range depends on the choice of ground-cost $c$. Second, it is only defined in the support of $\hat{P}$. We offer a solution to both of these problems.
@@
Concerning interpretability, we propose to transform it using the Cumulative Distribution Function (CDF) of its values. Let $P_{\mathcal{T}}$ be the probability measure associated with $\mathcal{T}(x)$. The CDF is simply $F_{\mathcal{T}}(t)=P_{\mathcal{T}}((-\infty,t))$. Naturally, since $P_{\mathcal{T}}$ is not available, it may be approximated from samples $\{\mathcal{T}_{i}\}_{i=1}^{n}$, obtained through equation 9. We do so through, Kernel Density Estimation (KDE),
$\hat{k}(t)$ $=\dfrac{1}{n\sigma}\sum_{i=1}^{n}\phi\biggr{(}\dfrac{t-\mathcal{T}_{i}}{\sigma}\biggr{)},$ (10)
@@
where $K$ is a kernel function (e.g., the Gaussian kernel $\phi(x)=\exp(-\nicefrac{{x^{2}}}{{2}})$), and $\sigma$ is the bandwidth parameter, controlling the smoothness of $\hat{k}$ and determined through Scott’s rule (Scott, 1979). Equation 10 provide an approximation for the density of transportation efforts $\{\mathcal{T}_{i}\}_{i=1}^{n}$. The CDF can be easily calculated from this estimate, via,
[…]
The CDF is appealing, as it is a monotonic function over transportation efforts $t\in\mathbb{R}$, and it takes values on $[0,1]$, both of which are desirable properties for an anomaly score.
@@
The issue on how to extrapolate the anomaly score for new samples remain. For instance, even if we use the CDF values as anomaly scores, one needs to recalculate $t=\hat{\mathcal{T}}(\mathbf{x})$ for a new sample, before evaluating $\hat{H}(t)$ or $\hat{K}(t)$. As we previously discussed, this is challenging, as $\hat{\mathcal{T}}$ is only defined on the support of $\hat{P}$. A naive approach would be to append $\mathbf{x}$ to the set $\{\mathbf{x}_{i}^{(P)}\}_{i=1}^{n}$ and solve a MROT problem again. Naturally, this is not feasible, as solving an OT problem for each new sample is computationally expensive.
@@
In this paper we put forth the more efficient idea of modeling the relationship $\mathbf{x}\mapsto\hat{K}(\hat{\mathcal{T}}(\mathbf{x}))$ from the samples we are given, that is, we create a labeled dataset $\{\mathbf{x}_{i}^{(P)},t_{i}\}_{i=1}^{n}$, where $t_{i}=\hat{\mathcal{T}}(\mathbf{x}_{i}^{(P)})$. Our anomaly score comes, then, through a function $\psi(\mathbf{x}_{i}^{(P)})$ fit to the labeled dataset through regression. Fitting $\psi$ can be done with standard regression tools, such as Ordinary Least Squares (OLS), Support Vector Regression (SVR) (Smola and Schölkopf, 2004), nearest neighbors, or gradient boosting (Friedman, 2002). In general, one can expect the relationship between $\mathbf{x}_{i}^{(P)}$ and $\hat{K}(\hat{\mathcal{T}}(\mathbf{x}_{i}^{(P)}))$ to be non-linear, hence, it is generally necessary to use a non-linear regression model. We show a summary of our strategy in Algorithm 1.
1 function mrot($\mathbf{X}^{(P_{S})},\epsilon,k$)
@@ 2       $\gamma_{k,\epsilon}^{\star}=\text{MROT}_{k,\epsilon}(\hat{P})$;
3       $t_{i}\leftarrow\sum_{j=1}^{n}(\nicefrac{{(\gamma_{k,\epsilon}^{\star})_{ij}}}{{p_{i}}})\tilde{c}_{k}(x_{i}^{(P)},x_{j}^{(P)})$;
4       $\hat{k}\leftarrow\text{KDE}(\{t_{i}\}_{i=1}^{n})$;
5       $\psi\leftarrow\text{Regression}(\{x_{i}^{(P)},\hat{K}(t_{i})\}_{i=1}^{n})$;
@@ 4 Experiments
We divide our experiments in 4 parts. Section 4.1 shows a toy example, going through all the steps in our algorithm. Section 4.2 shows our results on AdBench (Han et al., 2022). Section 4.3 shows our experiments in fault detection on the Tennessee Eastman Process (Montesuma et al., 2024b; Reinartz et al., 2021). Finally, Section 4.4 explores the robustness of our methods to various hyper-parameters and design choices.
4.1 An introductory toy example
Before diving into comparing our method with prior art, we give an introductory example that illustrates how we create anomaly scores out of samples. In this example, we sample normal examples $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},0.25\mathbf{I}_{2})$, and anomalous samples $\mathbf{y}_{j}\sim\mathcal{N}([-3,-3],0.01\mathbf{I}_{d})$, where $\mathbf{I}_{2}$ is a $2\times 2$ identity matrix. The dataset for this toy example consists of the concatenation $\{\mathbf{x}_{1},\cdots,\mathbf{x}_{n},\mathbf{y}_{1},\cdots,\mathbf{y}_{m}\}$, where $n=500$ and $m=25$, which means that roughly $5\%$ of the total number of samples are anomalies. In the following, we compare our engineered cost with a regularized Coulomb interaction, $C_{ij}=(1+\lVert\mathbf{x}_{i}^{(P)}-\mathbf{x}_{j}^{(P)}\rVert_{2})^{-1}$. This cost is shown alongside our proposed engineered cost in Figure 2. Note that, in Figures 2 (b - d) the lower-right corner of the matrices correspond to anomalies.
(a) Toy example.
(b) Euclidean cost.
(c) Engineered cost.
(d) Coulomb cost.
Figure 2: In (a), we show the samples for our toy example. In (b), (c) and (d), we show the ground-cost $C_{ij}$ between samples in (a), using the Euclidean distance, our engineered ground-cost (equation 6), and the regularized Coulomb interaction cost.
Figure 2 clearly illustrates that directly using the Euclidean cost results in a trivial OT solution, because the diagonal entries of $C_{ij}$ are zero. This is not the case for our engineered cost, and the regularized Coulomb interaction, shown in Figures 2 (c) and (d). To stress this idea, we show in Figure 3 the transportation plans acquired by MROT and using the regularized Coulomb cost. While both strategies achieve similar transportation matrices, in the second case the anomalies send and receive less mass than normal samples. This is somewhat expected, as the anomalous samples are clustered together in a tight region of $\mathbb{R}^{2}$, resulting in a high cost. As a result, these samples are encouraged to send their mass elsewhere.
@@ 4.1 An introductory toy example
Figure 3: In (a), we show the MROT transportation plan (left side) and strategy (right side). Likewise, in (b) we show the transportation plan and strategy for the regularized Coulomb cost. While both transportation strategies achieve similar results, in (b) the anomalies (lower right corner) sends and receives much less mass than other normal samples.
Here comes an important distinction between our approach and simply using repulsive costs in OT. Considering our last remark about the regularized Coulomb interaction, samples send their mass to distant regions of space, thus reducing $C_{ij}$. As a result, anomalies will incur in a smaller transportation cost than normal samples. Although counterintuitive, one can still construct a detection rule out of this idea, as we show below.
@@
As we discussed throughout section 3.3, the scores in Figure 4 are only defined in the support of $\hat{P}$. We then explore the regression of this score through regression, which can be done with a variety of standard regression algorithms. In Figure 5 we show the results for eXtreme Gradient Boosting (XGBoost) (Chen and Guestrin, 2016) and SVR (Smola and Schölkopf, 2004), which both define the anomaly score over the whole ambient space $\mathbb{R}^{2}$. As an important remark, the relationship between $x_{i}^{(P)}$ and its score is likely non-linear. Indeed, this idea is evidenced in Figure 5. As a result, one needs a non-linear regression algorithm.
@@
Figure 5: Regression of the anomaly score for MROT (ours, a) and OT with repulsive costs (b).
@@ 4.2 Comparison on AdBench
AdBench (Han et al., 2022) is a benchmark in AD with 57 different kinds of datasets, grouped into 47, 5 and 5 real-world, vision and Natural Language Processing (NLP) datasets. In our experiments, we focus on real-world and natural-language datasets. We compare, in total, 17 methods, grouped into classical, diffusion-based and our proposed methods based on OT. For classic methods, we consider Isolation Forest (IsoF) (Liu et al., 2008), One Class SVM (OCSVM) (Schölkopf et al., 1999), $k-$NN, LOF (Breunig et al., 2000), CBLOF (He et al., 2003), Empirical Cumulative Distribution Functions (ECOD) (Li et al., 2022), Copula-Based OD (COPOD) (Li et al., 2020), Lightweight On-line Detector of Anomalies (LODA) (Pevnỳ, 2016), Feature Bagging (Lazarevic and Kumar, 2005), and Histogram-based Outlier score (HBOS) (Goldstein and Dengel, 2012). For diffusion based, we consider the 4 variants of Diffusion Time Estimation (DTE) (Livernoche et al., 2024). Finally, we consider OT with repulsive costs Di Marino et al. (2017), and our proposed MROT.
@@
Figure 6: AdBench result summary. AUC-ROC per dataset is available in the appendix. We compare, in total, 17 algorithms over 47 real-world datasets (a) and 5 NLP datasets (b). For real-world datasets, our MROT has state-of-the-art performance, especially when compared with recently proposed flow-based models. For higher dimensional data, our method
@@
Even though our method has superior performance on real-world datasets, the NLP benchmarks reveal limitations of OT based AD. A first limitation of using MROT comes from the scalability of OT. As a linear program, it has at least $\mathcal{O}(n^{2})$ storage, and $\mathcal{O}(n^{3}\log n)$ computational complexity, where $n$ is the number of samples. In our experiments, we limited the number of samples to $n=20,000$, by down-sampling bigger datasets. As reported in Figure 6 (a), and our detailed results are given in the Appendix, this process does not affect performance. Furthermore, from Figure 6 (b), we see that our method struggles in high dimensional anomaly detection, such as those in the NLP datasets of Han et al. (2022). This limitation stems from the use of OT in high-dimensions (Montesuma et al., 2025, Section 8.1).
@@ 4.3 Tennessee Eastman Process
In this section, we focus in comparing anomaly detection methods for fault detection. To that end, we use the Tennessee Eastmann (TE) process benchmark (Downs and Vogel, 1993), especially the simulations of Reinartz et al. (2021), pre-processed by Montesuma et al. (2024b). This benchmark is composed by simulations of a large-scale chemical plant. Besides the normal state, there are 28 faults associated with different parts of the plant. An interesting feature of this benchmark is that there are different modes of operation. Ideally, an anomaly detection algorithm trained in one mode should generalize to other modes, but this is not guaranteed due to distributional shift. Such challenge is the subject of domain adaptation (Montesuma et al., 2023, 2024b, 2024c, 2025), which is beyond the scope of this work. We refer readers to the aforementioned papers for further details.
From the original simulations in Reinartz et al. (2021), we consider $100$ simulations for the normal state. These simulations last for $600$ hours, with a time-step of $1$ hour, and concern $34$ sensors measuring different physical and chemical properties. On top of these $100$ simulations, we take $1$ faulty simulation for each kind of fault, leading to a total of $128$ simulations. Based on this set of simulations, we extract windows of $20$ hours from the original signals. Each window is then considered a sample for anomaly detection. On each window, we compute the mean, and standard deviation of each variable, leading to vectors $\mu,\sigma\in\mathbb{R}^{34}$. We use the concatenation of these vectors as features for anomaly detection. These steps lead to $6$ datasets with $3840$ samples, $840$ ($21.875\%$) of them being anomalous, one for each mode of operation. We thus divide our discussion into two parts: fault detection, and cross-domain fault detection.
@@
In Figure 7, we visualize the Tennessee Eastman data using t-SNE. We scatter the embeddings of each mode’s data, showing that most faulty samples cluster outside the normal data cluster. However, some faulty samples are close to normal ones. This phenomenon is expected, as the effect of faults evolves over time. As a result, windows taken in early stages of simulation resemble those of normal samples.
@@
4.4 Robustness

### v2

@@ 3.4 Building and Generalizing an Anomaly Score
Through MROT, we want to build a score for the samples based on how anomalous they are. Assuming that anomalies lie in a low density region of space, our MROT problem forces those samples to send their mass to distant parts of the feature space. In contrast, normal samples can send their mass to the immediate neighborhood outside the exclusion zone. As a result, we can sort out normal from anomalous samples using the transportation effort,
@@
$t_{i}=\mathcal{T}(x_{i}^{(P)})=\sum_{j=1}^{n}\dfrac{\gamma_{ij}^{\star}}{p_{i}}\tilde{c}_{k}(x_{i}^{(P)},x_{j}^{(P)}),$ (10)
@@
where $p_{i}$ is the importance of the $i-$th sample (e.g, $p_{i}=n^{-1}$ for uniform importance). Interestingly, equation 10 is similar to the barycentric map, widely used in domain adaptation Courty et al. (2016). $\mathcal{T}$ has 2 shortcomings as an anomaly score. First, it is hardly interpretable, as its range depends on the choice of ground-cost $c$. Second, it is only defined in the support of $\hat{P}$. We offer a solution to both of these problems.
@@
Concerning interpretability, we propose to transform it using the Cumulative Distribution Function (CDF) of its values. Let $P_{\mathcal{T}}$ be the probability distribution associated with $\mathcal{T}(x)$. The CDF is simply $F_{\mathcal{T}}(t)=P_{\mathcal{T}}((-\infty,t))$. Naturally, since $P_{\mathcal{T}}$ is not available, it may be approximated from samples $\{t_{i}\}_{i=1}^{n}$, obtained through equation 10. We do so through, Kernel Density Estimation (KDE),
$\hat{P}_{\mathcal{T}}(t)$ $=\dfrac{1}{n\sigma}\sum_{i=1}^{n}\phi\biggr{(}\dfrac{t-t_{i}}{\sigma}\biggr{)}\text{, and, }\hat{F}_{\mathcal{T}}(t)=\int_{-\infty}^{t}\hat{P}_{\mathcal{T}}(s)ds$ (11)
@@
where $\phi$ is a kernel function (e.g., the Gaussian kernel $\phi(x)=\exp(-\nicefrac{{x^{2}}}{{2}})$), and $\sigma$ is the bandwidth, controlling the smoothness of $\hat{P}_{\mathcal{T}}$ and determined through Scott’s rule (Scott, 1979). Equation 11 is an approximation for the density of transportation efforts $\{t_{i}\}_{i=1}^{n}$. The CDF is appealing, as it is a monotonic function over transportation effort $t\in\mathbb{R}$, and it takes values on $[0,1]$, both of which are desirable for an anomaly score.
@@
The question of how to extrapolate the anomaly score for new samples remains. For example, even if we use the CDF values as anomaly scores, we need to recalculate $t=\hat{\mathcal{T}}(\mathbf{x})$ for a new sample $\mathbf{x}\sim P$, which is challenging, as $\hat{\mathcal{T}}$ is only defined on the support of $\hat{P}$. A naive approach would be to append $\mathbf{x}$ to the set $\{\mathbf{x}_{i}^{(P)}\}_{i=1}^{n}$ and solve a MROT problem again. Naturally, this is not feasible, as solving an OT problem for each new sample is computationally expensive.
@@
In this paper, we present the more efficient idea of modeling the relationship $\mathbf{x}\mapsto\hat{F}_{\mathcal{T}}(\hat{\mathcal{T}}(\mathbf{x}))$ from the samples we receive, that is, we create a labeled data set $\{\mathbf{x}_{i}^{(P)},t_{i}\}_{i=1}^{n}$, where $t_{i}=\hat{\mathcal{T}}(\mathbf{x}_{i}^{(P)})$. Our anomaly score comes, then, through a function $\psi(\mathbf{x}_{i}^{(P)})$ fit to the labeled dataset through regression. The fitting $\psi$ can be done with standard regression tools, such as Ordinary Least Squares (OLS), Support Vector Regression (SVR) (Smola and Schölkopf, 2004), nearest neighbors or gradient boosting (Friedman, 2002). In general, one can expect the relationship between $\mathbf{x}_{i}^{(P)}$ and $\hat{F}_{\mathcal{T}}(\hat{\mathcal{T}}(\mathbf{x}_{i}^{(P)}))$ to be non-linear, hence, it is generally necessary to use a non-linear regression model. We show a summary of our strategy in Algorithm 1.
1 function mrot($\mathbf{X}^{(P)},\epsilon$)
@@ 2       $\gamma_{k,\epsilon}^{\star}=\text{MROT}_{k,\epsilon}(\hat{P})$;
3       $t_{i}\leftarrow\sum_{j=1}^{n}(\nicefrac{{(\gamma_{\epsilon}^{\star})_{ij}}}{{p_{i}}})\tilde{c}(\mathbf{x}_{i}^{(P)},\mathbf{x}_{j}^{(P)})$;
4       $\hat{P}_{\mathcal{T}}\leftarrow\text{KDE}(\{t_{i}\}_{i=1}^{n})$;
5       $\psi\leftarrow\text{Regression}(\{\mathbf{x}_{i}^{(P)},\hat{F}_{\mathcal{T}}(t_{i})\}_{i=1}^{n})$;
@@ 4 Experiments
We divide our experiments in 3 parts. Section 4.1 shows our results on AdBench (Han et al., 2022). Section 4.2 shows our experiments in fault detection on the Tennessee Eastman Process (Montesuma et al., 2024b; Reinartz et al., 2021). Finally, Section 4.3 explores the robustness of our methods to various hyper-parameters and design choices.
In the following, we consider adaptive neighborhoods $\mathcal{N}_{k}(\mathbf{x})$ which take the $k-$nearest neighborhood of $\mathbf{x}$, and set $C_{ij}=\text{max}_{\ell=1,\cdots,n}C_{i\ell}$, where $j\in\mathcal{N}_{k}(\mathbf{x}_{i}^{(P)})$. Before going through our experiments, we show a toy example, going through all the steps in our algorithm. See Figure 3 for a comparison.
An introductory example. Before comparing our method with prior art, we give an introductory example that illustrates how we create anomaly scores out of samples. In this example, we sample normal examples $\mathbf{x}_{i}\sim\mathcal{N}(\mathbf{0},0.25\mathbf{I}_{2})$, and anomalous samples $\mathbf{y}_{j}\sim\mathcal{N}([-3,-3],0.01\mathbf{I}_{2})$, where $\mathbf{I}_{2}$ is a $2\times 2$ identity matrix. The dataset for this toy example consists of the concatenation $\mathbf{X}^{(P)}=\{\mathbf{x}_{1},\cdots,\mathbf{x}_{n},\mathbf{y}_{1},\cdots,\mathbf{y}_{m}\}$, where $n=500$ and $m=25$, which means that roughly $5\%$ of the total number of samples are anomalies. These samples are shown in Figure 5.
Figure 4: Toy example.
We compare our engineered cost with a regularized Coulomb interaction, $C_{ij}=(1+\lVert\mathbf{x}_{i}^{(P)}-\mathbf{x}_{j}^{(P)}\rVert_{2})^{-1}$. This cost is shown alongside our proposed engineered cost in Figure 4. Note that, in Figures 5 (a - c) the lower right corner of the matrices corresponds to anomalies.
From Figure 5, OT with the Euclidean cost results in a trivial OT solution, because the diagonal entries $C_{ij}$ are zero. Hence, $\gamma^{\star}=\mathbf{I}_{n}$ achieves 00 transportation cost. This is not the case for our engineered cost, and the regularized Coulomb interaction, shown in Figures 5 (c) and (d). To emphasize this idea, we show in Figure 6 the transportation plans acquired by MROT, and the regularized Coulomb cost. For the Coulomb cost, the anomalies send and receive less mass than normal samples. This is somewhat expected, as the anomalous samples are clustered together in a tight region of $\mathbb{R}^{2}$, resulting in a high cost. As a result, these samples are encouraged to send their mass elsewhere. This phenomenon does not happen with our engineered cost, which encourages samples to send their mass just outside to their exclusion zone.
(a) Euclidean cost.
(b) Engineered cost.
(c) Coulomb cost.
Figure 5: In (a), (b) and (c), we show the ground-cost $C_{ij}$ between samples in Figure 4, using the Euclidean distance, our engineered ground-cost (equation 6), and the regularized Coulomb interaction cost.
@@
Figure 6: In (a), we show the MROT transportation plan (left side), the matching (center) and the transport plan (right side). Likewise, in (b) we show the transportation plan and strategy for the regularized Coulomb cost. As shown in (b) left side and right side, the anomalies (lower right corner) sends and receives much less mass than other normal samples.
Hence, there is an important distinction between our approach and using repulsive costs in OT. Considering our last remark about the regularized Coulomb interaction, samples send their mass to distant regions of space, thus reducing $C_{ij}$. As a result, anomalies will incur in a smaller transportation cost than normal samples. Although counterintuitive, one can still construct a detection rule out of this idea.
@@
As we discussed throughout section 3.4, the scores in Figure 7 are only defined in the support of $\hat{P}$. We then explore the regression of this score through regression, which can be done with a variety of standard regression algorithms. In Figure 8 we show the results for eXtreme Gradient Boosting (XGBoost) (Chen and Guestrin, 2016) and SVR (Smola and Schölkopf, 2004), which both define the anomaly score over the whole ambient space $\mathbb{R}^{2}$. As an important remark, the relationship between $\mathbf{x}_{i}^{(P)}$ and its score is likely non-linear. Indeed, this idea is evidenced in Figure 8. As a result, one needs a non-linear regression algorithm.
@@
Figure 8: Regression of the anomaly score for MROT (ours, a) and OT with repulsive costs (b). While our method attributes higher scores to anomalous regions of the space, OT with repulsive costs does the inverse.
@@ 4.1 Comparison on AdBench
AdBench (Han et al., 2022) is a benchmark in AD with 57 different kinds of datasets, grouped into 47, 5 and 5 real-world, vision and Natural Language Processing (NLP) datasets. In our experiments, we focus on real-world and natural-language datasets. We compare, in total, 17 methods, grouped into classical, diffusion-based and our proposed methods based on OT. For classic methods, we consider Isolation Forest (IsoF) (Liu et al., 2008), OCSVM (Schölkopf et al., 1999), $k-$NN, LOF (Breunig et al., 2000), CBLOF (He et al., 2003), Empirical Cumulative Distribution Functions (ECOD) (Li et al., 2022), Copula-Based OD (COPOD) (Li et al., 2020), Lightweight On-line Detector of Anomalies (LODA) (Pevnỳ, 2016), Feature Bagging (Lazarevic and Kumar, 2005), and Histogram-based Outlier score (HBOS) (Goldstein and Dengel, 2012). For diffusion based, we consider the 4 variants of Diffusion Time Estimation (DTE) (Livernoche et al., 2024). We also consider OT with repulsive costs Di Marino et al. (2017), and MROT (ours).
@@
Figure 9: AdBench result summary. AUC-ROC and AUC-PR per dataset is available in the appendix. We compare, in total, 17 algorithms over 47 real-world datasets (a) and 5 NLP datasets (b). For real-world datasets, our MROT has state-of-the-art performance, especially when compared with recently proposed flow-based models. For higher dimensional data, our method
@@
Here, we highlight two limitations of our method. The first limitation of using MROT comes from the scalability of OT. As a linear program, it has at least $\mathcal{O}(n^{2})$ storage, and $\mathcal{O}(n^{3}\log n)$ computational complexity, where $n$ is the number of samples. In our experiments, we limited the number of samples to $n=20,000$, by down-sampling larger datasets. As reported in Figure 9 (a), and our detailed results given in the Appendix, this process does not affect performance. Furthermore, from Figure 9 (b), we see that our method struggles in high dimensional AD, such as those in the NLP datasets of Han et al. (2022). This limitation stems from the use of OT in high-dimensions (Montesuma et al., 2024a, Section 8.1).
@@ 4.2 Tennessee Eastman Process
In this section, we focus in comparing anomaly detection methods for fault detection in control systems. In this context, as defined by Isermann (2006), a fault is an anomaly in at least one of the system’s variables. To that end, we use the Tennessee Eastmann (TE) process benchmark (Downs and Vogel, 1993), especially the simulations of Reinartz et al. (2021), pre-processed by Montesuma et al. (2024b).
This benchmark is composed by simulations of a large-scale chemical plant, from which a collection of 34 physical and chemical quantities are measured over time. There are a total of 29 states for this plant: 1 healthy state, and 28 faulty states. Furthermore, the plant can operate under 6 different modes of operation, depending on the specified production requirements. We refer readers to the aforementioned references, as well as our appendix, for detailed descriptions of the variables, faults and modes of operation.
@@
From the original simulations in Reinartz et al. (2021), we consider $100$ simulations for the normal state. These simulations last for $600$ hours, with a time-step of $1$ hour, and concern $34$ sensors measuring different physical and chemical properties. On top of these $100$ simulations, we take $1$ faulty simulation for each kind of fault, leading to a total of $128$ simulations. Based on this set of simulations, we extract windows of $20$ hours from the original signals. Each window is then considered a sample for anomaly detection. On each window, we compute the mean, and standard deviation of each variable, leading to vectors $\mu,\sigma\in\mathbb{R}^{34}$. We use the concatenation of these vectors as features for anomaly detection. These steps lead to $6$ datasets with $3840$ samples, $840$ ($21.875\%$) of them being anomalous, one for each mode of operation.
@@
We start our analysis by embedding the obtained features in $\mathbb{R}^{2}$ through t-Stochastic Neighbor Embeddings (TSNE) (Van der Maaten and Hinton, 2008), which is shown in Figure 10. We scatter the embeddings of each mode’s data, showing that most faulty samples cluster outside the normal data cluster. However, some faulty samples are close to normal ones. This phenomenon is expected, as the effect of faults evolves over time. As a result, windows taken in early stages of simulation resemble those of normal samples.
4.3 Ablations
Hyper-parameter robustness. Here, we analyze the robustness of our method with respect to the entropic regularization penalty $\epsilon$, and the number of nearest neighbors $k$ in $\mathcal{N}_{k}$. In our experiments, we evaluated our method on the values $\epsilon\in\{0,10^{-2},10^{-1},10^{0}\}$, where $\epsilon=0$ implies the use of exact OT, that is, linear programming. For MROT, we use $k\in\{5,10,20,\cdots,50\}$. Note that while $\epsilon$ is related to transportation plan, $k$ is linked to the ground-cost. We summarize our results in Figure 12, where we present box-plots over the AUC-ROC scores on each dataset in the AdBench benchmark, for each combination of hyper-parameters.